# An algebraic theory of Łojasiewicz exponents

### v1

@@ 10.7. Monomial filtrations in the toric setting
(1) By Theorem 5.6 we have the universal bound $\mathcal{L}_{\mathfrak{b}_{\bullet}}(\mathfrak{a}_{\bullet})\geq\sup_{v}\frac{v(\mathfrak{a}_{\bullet})}{v(\mathfrak{b}_{\bullet})}$ over divisorial valuations $v$ centered at $\mathfrak{m}_{S}$. Restricting to toric divisorial valuations $v=v_{u}$ gives the first inequality. The expression in terms of $\operatorname{NP}(\mathfrak{a}_{\bullet})$ and $\operatorname{NP}(\mathfrak{b}_{\bullet})$ follows from Lemma 10.14.
(2) This is exactly Lemma 10.15 applied to the functions $u\mapsto v_{u}(\mathfrak{a}_{\bullet})$ and $u\mapsto v_{u}(\mathfrak{b}_{\bullet})$.
(3) Under the stated testing equivalence, the finite testing hypothesis (H1) of Theorem 5.8 holds with $V=\{v_{u}:u\in\mathrm{Ray}(\Sigma)\}$, hence Theorem 5.8 yields the desired equality. ∎
@@ 11.2. Newton nondegeneracy and toric compactification
We will use a toric compactification adapted to the Newton polyhedron at infinity following Khovanskii [33]. Let $\Sigma_{\Gamma}$ denote the normal fan of $\Gamma(F)$ and let $\Sigma_{\mathbb{P}^{n}}$ denote the standard fan of $\mathbb{P}^{n}$. Choose a smooth complete fan $\Sigma$ refining both $\Sigma_{\Gamma}$ and $\Sigma_{\mathbb{P}^{n}}$ (such a refinement exists by toric resolution of singularities; see [33, 42], [11, Chapter 11] and [18, Chapter 2]). Then, the associated toric variety $X_{\Sigma}$ is a smooth projective toric compactification of the torus $T=(k^{*})^{n}$, and the refinement $\Sigma\succ\Sigma_{\mathbb{P}^{n}}$ induces a proper toric morphism $\pi_{\Sigma}:X_{\Sigma}\longrightarrow\mathbb{P}^{n}$ extending the standard open embedding $T\hookrightarrow\mathbb{P}^{n}$ (see [11, Chapter 3] and [18, Chapters 1–2]). Moreover, since $\Sigma$ also refines the normal fan of $\Gamma(F)$, the toric boundary divisors of $X_{\Sigma}$ correspond to weight vectors governing the face systems (initial forms) of $F$, which control the asymptotic behavior at infinity as in Khovanskii’s compactification.
@@
We work on a toric compactification attached to $\Gamma(F)$ and then descend to divisorial valuations of $R_{\xi}$.
@@
Step 2: boundary intersections of the graph and Newton nondegeneracy. Let $\Gamma_{F}^{\Sigma}\subseteq X_{\Sigma}\times\mathbb{P}^{r}$ be as before, and let $\mu_{\Sigma}:\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$ be the induced proper birational morphism.
Fix a torus–invariant divisor $E\subseteq D_{\operatorname{Tor}}$ and consider the locally closed stratum $O(E)\subseteq X_{\Sigma}$. Work on the target chart $\{W_{0}\neq 0\}\subseteq\mathbb{P}^{r}$, with affine coordinates $y_{j}:=W_{j}/W_{0}$. Restricting to $O(E)\times\mathbb{A}^{r}$, we work in an affine toric chart $U_{\sigma}=\operatorname{Spec}k[\sigma^{\vee}\cap M]$ containing $O(E)$, where $\sigma\in\Sigma$ is a cone whose rays include $w_{E}$. In the corresponding toric coordinates $u_{1},\dots,u_{n}$, each coordinate $x_{i}$ is a monomial in the $u_{j}$, and each polynomial $f_{j}(x)$ admits a decomposition of the form
$f_{j}(x)=u^{m_{j}}\Bigl((f_{j})_{w_{E}}(u)+\sum_{\langle w_{E},\alpha\rangle>m_{j}}c_{j,\alpha}u^{\alpha}\Bigr),$
where $m_{j}=\min_{\alpha\in\Gamma(f_{j})}\langle w_{E},\alpha\rangle$, and $(f_{j})_{w_{E}}$ is the face polynomial consisting of those terms achieving this minimum. Consequently, the strict transform of the graph in $U_{\sigma}\times\mathbb{A}^{r}$ meets $O(E)\times\mathbb{A}^{r}$ if and only if the face system
$(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0$
admits a solution in the torus $O(E)\cong(k^{*})^{n}$. This is the standard initial-form criterion for strict transforms in toric modifications; see, for example, [33, 42]. In particular,
(11.7) $\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing\Leftrightarrow(f_{1})_{w_{E}}=\cdots=(f_{r})_{w_{E}}=0\ \text{has no solution in }(k^{*})^{n}.$
Since $F$ is Newton nondegenerate at infinity, the right-hand condition holds for every $w_{E}\in\mathbb{R}_{>0}^{n}$, and so (11.7) implies: for every boundary divisor $E$ with $w_{E}\in\mathbb{R}_{>0}^{n}$,
$\Gamma_{F}^{\Sigma}\cap\bigl(O(E)\times\{W_{0}\neq 0\}\bigr)=\varnothing,$
i.e. every point of $\Gamma_{F}^{\Sigma}$ lying over $O(E)$ maps into the hyperplane at infinity $H_{Y}=\{W_{0}=0\}$. Equivalently, $\rho_{\Sigma}^{-1}(H_{Y})$ contains $\mu_{\Sigma}^{-1}(\pi_{\Sigma}^{-1}(H_{X}))$ along all such boundary strata.
Step 3: descend to divisorial valuations over $\xi$. Let $v$ be a divisorial valuation of the local domain $R_{\xi}=\mathcal{O}_{\overline{\Gamma}_{F},\xi}$. Choose a normal modification $\nu:W\to\overline{\Gamma}_{F}$ and a prime divisor $E^{\prime}\subseteq W$ with center $\xi$ such that $v=\operatorname{\mathrm{ord}}_{E^{\prime}}$ on $R_{\xi}$. Take a common normal refinement $W^{\prime}\to W$ and $W^{\prime}\to\Gamma_{F}^{\Sigma}$ dominating both models, and let $\widetilde{E}^{\prime}\subseteq W^{\prime}$ be the strict transform of $E^{\prime}$. Then
@@
Assume $v(\mathscr{I}_{X,\xi})>0$. This means that $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, so its image in $X_{\Sigma}$ is contained in the toric boundary lying over $H_{X}$. Hence, $\widetilde{E}^{\prime}$ dominates (after possibly replacing $\Sigma$ by a further refinement) a torus–invariant boundary divisor $E\subseteq D_{\operatorname{Tor}}$ with ray vector $w_{E}\in\mathbb{R}_{>0}^{n}$. Indeed, divisors corresponding to rays lying on coordinate faces of the cone $\mathbb{R}_{\geq 0}^{n}$ map into the coordinate boundary of $\mathbb{A}^{n}_{k}\subseteq\mathbb{P}^{n}$ rather than into the hyperplane at infinity $H_{X}=\{X_{0}=0\}$. Since $\widetilde{E}^{\prime}$ lies over $D_{X}=\pi^{-1}(H_{X})$, it must dominate a torus-invariant divisor corresponding to a ray in the relative interior $\mathbb{R}_{>0}^{n}$ of the cone. This is a standard property of toric compactifications attached to Newton polyhedra; see [33, 42]. By Step 2, $\Gamma_{F}^{\Sigma}$ over $O(E)$ maps into $H_{Y}$, so $\widetilde{E}^{\prime}$ is contained in the pullback of $D_{Y}=\rho^{-1}(H_{Y})$. Equivalently, $\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\mathscr{I}_{Y}\mathcal{O}_{W^{\prime}})>0$, i.e. $v(\mathscr{I}_{Y,\xi})>0$.
This proves $v(\mathscr{I}_{X,\xi})>0\Rightarrow v(\mathscr{I}_{Y,\xi})>0$. The second statement is immediate. ∎
@@ 11.3. Newton nondegenerate at infinity and Łojasiewicz exponent
    for every $\xi\in D_{X}\cap D_{Y}$, after base change to $\operatorname{Spec}(R_{\xi})$ the ideals $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
@@
For (2), fix $\xi\in D_{X}\cap D_{Y}$ and base change $Z\to\overline{\Gamma}_{F}$ to $\operatorname{Spec}(R_{\xi})$. On the smooth scheme $Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$, the pullbacks of $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ define Cartier divisors supported on the pullback of $(D_{X})_{\text{red}}+(D_{Y})_{\text{red}}$ and, hence, on the strict transforms of those $E_{i}$ meeting $\mu^{-1}(\xi)$. In particular, both ideals become invertible on $Z_{\xi}$ with support contained in those components, as claimed. ∎
@@ 12.2. Uniform verification on a fixed toric model
$\mathfrak{a}_{t}\ =\ \pi_{*}\mathcal{O}_{Y}\Big(-\sum_{i=1}^{N}\Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\,E_{i}\Big).$
In particular, for any other family $\mathfrak{b}_{\bullet}$} of the same form (possibly with different coefficients), the containment relations $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ are determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E_{i}}\}_{i=1}^{N}$, i.e., the hypotheses of the finite-testing criterion in Theorem 5.8 are satisfied with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$.
@@
[…]
@@
$\Big(\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\Big)\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}\big(-D(t)\big).$
@@
$\mathfrak{a}_{t}\ =\ \overline{\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}}\ =\ \pi_{*}\mathcal{O}_{Y}\big(-D(t)\big).$
Step 2: replace $\lceil\lambda_{j}t\rceil$ by linear bounds in $t$. Write $D(t)=\sum_{i}d_{i}(t)E_{i}$ as above. Since $d_{i}(t)$ is an integer-valued, piecewise linear function in $t$ with slopes $\sum_{j}\lambda_{j}a_{ij}$, we have
$d_{i}(t)\ =\ \Big\lfloor t\cdot\sum_{j=1}^{r}\lambda_{j}a_{ij}\Big\rfloor\ +\ \epsilon_{i}(t),\quad\epsilon_{i}(t)\in\{0,1,\dots,r\}.$
Absorbing the bounded error into flooring yields the displayed formula; this is the same bookkeeping already used in Section 5 when passing between integer- and real-indexed families.
Step 3: finite testing. For any $p,q$, the condition $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ is equivalent to the inequality of divisors on $Y$: $D_{\mathfrak{b}_{\bullet}}(q)\ \geq\ D_{\mathfrak{a}_{\bullet}}(p)$, i.e., coefficientwise inequalities along $E_{i}$. Equivalently, $\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{q})\geq\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{p})$ for all $i$. This is precisely the finite-testing hypothesis in Theorem 5.8 with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$. ∎

### v2

@@ 10.7. Monomial filtrations in the toric setting
By Theorem 5.6, we have
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\geq\sup\limits_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}.$
By Corollary 10.16, the latter supremum is equal to $\max\limits_{\{j:\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})>0\}}\frac{\widetilde{v}_{u_{j}}({\mathfrak{a}_{\bullet}})}{\widetilde{v}_{u_{j}}({\mathfrak{b}_{\bullet}})}.$ Thus, it remains to prove the reverse inequality.
Since $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is finitely generated, Theorem 5.9 gives a finite set of divisorial valuations $\mathcal{V}({\mathfrak{a}_{\bullet}})=\{v_{1},\dots,v_{r}\}$ such that
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})=\max_{1\leq i\leq r}\frac{v_{i}({\mathfrak{a}_{\bullet}})}{v_{i}({\mathfrak{b}_{\bullet}})},$
provided that $v_{i}({\mathfrak{b}_{\bullet}})>0$ for all $i$. We shall verify this positivity condition.
Indeed, since $\mathcal{R}({\mathfrak{a}_{\bullet}})$ is Noetherian, there exists $d\geq 1$ such that the $d$-th Veronese subalgebra $\mathcal{R}({\mathfrak{a}_{\bullet}})^{(d)}$ is standard graded. Equivalently, $\mathcal{R}({\mathfrak{a}_{\bullet}})^{(d)}$ is the ordinary Rees algebra of the monomial ideal $\mathfrak{a}_{d}$. Hence, the valuations in $\mathcal{V}({\mathfrak{a}_{\bullet}})$ are exactly the Rees valuations of $\mathfrak{a}_{d}$. By Theorem 10.11, these are toric divisorial valuations. Thus, there exist primitive lattice vectors $w_{1},\dots,w_{r}\in\sigma^{\circ}\cap N$ such that $v_{i}=v_{w_{i}}$ for all $i$.
Since $\mathfrak{b}_{\bullet}$} is linearly bounded with respect to $\mathfrak{q}$, there exists $c>0$ such that $v_{u}({\mathfrak{b}_{\bullet}})\geq c\,h_{\mathfrak{q}}(u)$ for all $u\in\sigma^{\circ}$. Therefore, $v_{w_{i}}({\mathfrak{b}_{\bullet}})\geq c\,h_{\mathfrak{q}}(w_{i})>0$ for every $i$, verifying the desired positive condition.
Now, each $w_{i}$ lies in $\sigma^{\circ}\cap N$, so $\frac{v_{w_{i}}({\mathfrak{a}_{\bullet}})}{v_{w_{i}}({\mathfrak{b}_{\bullet}})}\leq\sup\limits_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}$ for every $i$. Taking the maximum over $i$, we obtain the reverse inequality
$\mathcal{L}_{{\mathfrak{b}_{\bullet}}}({\mathfrak{a}_{\bullet}})\leq\sup_{u\in\sigma^{\circ}\cap N}\frac{v_{u}({\mathfrak{a}_{\bullet}})}{v_{u}({\mathfrak{b}_{\bullet}})}.$
The theorem is proved. ∎
@@ 11.2. Newton nondegeneracy and toric compactification
We will use a toric compactification adapted to the Newton polyhedron at infinity following Khovanskii [37]. Let $\Sigma_{\Gamma}$ denote the normal fan of $\Gamma(F)$ and let $\Sigma_{\mathbb{P}^{n}}$ denote the standard fan of $\mathbb{P}^{n}$. Choose a smooth complete fan $\Sigma$ refining both $\Sigma_{\Gamma}$ and $\Sigma_{\mathbb{P}^{n}}$ (such a refinement exists by toric resolution of singularities; see [37, 46], [14, Chapter 11] and [22, Chapter 2]). Then, the associated toric variety $X_{\Sigma}$ is a smooth projective toric compactification of the torus $T=(k^{*})^{n}$, and the refinement $\Sigma\succ\Sigma_{\mathbb{P}^{n}}$ induces a proper toric morphism $\pi_{\Sigma}:X_{\Sigma}\longrightarrow\mathbb{P}^{n}$ extending the standard open embedding $T\hookrightarrow\mathbb{P}^{n}$ (see [14, Chapter 3] and [22, Chapters 1–2]). Moreover, since $\Sigma$ also refines the normal fan of $\Gamma(F)$, the toric boundary divisors of $X_{\Sigma}$ correspond to weight vectors governing the face systems (initial forms) of $F$, which control the asymptotic behavior at infinity as in Khovanskii’s compactification.
@@
We will work with the notations in Definition 11.4 and Section 4.3.
@@
Step 2: boundary intersections of the graph and Newton nondegeneracy. Let $\Gamma_{F}^{\Sigma}\subseteq X_{\Sigma}\times\mathbb{P}^{r}$ be as before, and let $\mu_{\Sigma}:\Gamma_{F}^{\Sigma}\to\overline{\Gamma}_{F}$ be the induced proper birational morphism. We first claim that near $\mu_{\Sigma}^{-1}(\xi)$, one has
(11.7) $\operatorname{\mathrm{Supp}}(\mathscr{I}_{X}\mathcal{O}_{\Gamma_{F}^{\Sigma}})\subseteq\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$
Indeed, consider any point $z\in\operatorname{\mathrm{Supp}}(\mathscr{I}\mathcal{O}_{\Gamma_{F}^{\Sigma}})$. Then, $\mu_{\Sigma}(z)\in D_{X}$, so the image of $z$ in $X_{\Sigma}$ lies over the hyperplane at infinity $H_{X}$. Let $O(\sigma)\subseteq X_{\Sigma}$ be the torus orbit containing the image of $z$, where $\sigma\in\Sigma$. Since $O(\sigma)$ maps into $H_{X}$, the cone $\sigma$ is not contained in any coordinate face of $\mathbb{R}_{\geq 0}^{n}$. Equivalently, $\text{relint}(\sigma)\cap\mathbb{R}_{>0}^{n}\neq\varnothing.$ Choose $w\in\text{relint}(\sigma)\cap\mathbb{R}_{>0}^{n}.$
We shall work in the affine toric chart $U_{\sigma}=\operatorname{Spec}k[\sigma^{\vee}\cap M]\subseteq X_{\Sigma}$ containing $O(\sigma)$. In toric coordinates on $U_{\sigma}$, each polynomial $f_{j}$ has an expansion of the form
$f_{j}=u^{m_{j}}\Bigl((f_{j})_{w}(u)+\sum_{\langle w,\alpha\rangle>m_{j}}c_{j,\alpha}u^{\alpha}\Bigr),$
where $m_{j}=m_{w}(f_{j})=\min_{\alpha\in\Gamma(f_{j})}\langle w,\alpha\rangle$ and $(f_{j})_{w}$ is the face polynomial consisting of those terms achieving this minimum. By the standard initial-form criterion for strict transforms in toric compactifications (see, for example, [37, 46]), the graph closure and its normalization in $U_{\sigma}\times\mathbb{A}^{r}$ meet $O(\sigma)\times\{W_{0}\neq 0\}$ if and only if the face system
$(f_{1})_{w}=\cdots=(f_{r})_{w}=0$
has a solution in the torus $O(\sigma)\cong(k^{*})^{\,n-\dim\sigma}.$ Since $w\in\mathbb{R}_{>0}^{n}$ and $F$ is Newton nondegenerate at infinity, this system has no solution. Therefore, we conclude that
$\Gamma_{F}^{\Sigma}\cap\bigl(O(\sigma)\times\{W_{0}\neq 0\}\bigr)=\varnothing.$
@@
$z\in\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$
Since $z$ was arbitrary, this proves (11.7) near $\mu_{\Sigma}^{-1}(\xi)$.
Step 3: descend to divisorial valuations over $\xi$. Let $v$ be a divisorial valuation of $R_{\xi}$ with $v(\mathscr{I}_{X,\xi})>0.$ Choose a normal modification $\nu:W\to\overline{\Gamma}_{F}$ and a prime divisor $E^{\prime}\subseteq W$ with center $\xi$ such that $v=\operatorname{\mathrm{ord}}_{E^{\prime}}$ on $R_{\xi}$. Take a common normal refinement $\eta:W^{\prime}\to W$ and $\tau:W^{\prime}\to\Gamma_{F}^{\Sigma}$ dominating both models, and let $\widetilde{E}^{\prime}\subseteq W^{\prime}$ be the strict transform of $E^{\prime}$.
Since $\mathscr{I}_{X}$ is the ideal sheaf of the Cartier divisor $D_{X}$ on $\overline{\Gamma}_{F}$, its pullback to $W^{\prime}$ is invertible, and
$\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{X})=\operatorname{\mathrm{ord}}_{E^{\prime}}(\mathscr{I}_{X,\xi})=v(\mathscr{I}_{X,\xi})>0.$
Therefore, the generic point of $\widetilde{E}^{\prime}$ maps into $\operatorname{\mathrm{Supp}}(\mathscr{I}_{X}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ By (11.7), near $\mu_{\Sigma}^{-1}(\xi)$ this support is contained in $\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ Hence, the generic point of $\widetilde{E}^{\prime}$ also maps into $\operatorname{\mathrm{Supp}}(\mathscr{I}_{Y}\mathcal{O}_{\Gamma_{F}^{\Sigma}}).$ This implies that $\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{Y})>0.$
Moreover, $\mathscr{I}_{Y}$ is also the ideal sheaf of a Cartier divisor on $\overline{\Gamma}_{F}$. Thus,
$\operatorname{\mathrm{ord}}_{\widetilde{E}^{\prime}}(\tau^{*}\mathscr{I}_{Y})=\operatorname{\mathrm{ord}}_{E^{\prime}}(\mathscr{I}_{Y,\xi})=v(\mathscr{I}_{Y,\xi}).$
It follows that $v(\mathscr{I}_{Y,\xi})>0$, as desired. The second statement of the lemma is immediate. ∎
@@ 11.3. Newton nondegenerate at infinity and Łojasiewicz exponent
    for every $\xi\in D_{X}\cap D_{Y}$, after base change to $\operatorname{Spec}(R_{\xi})$ the ideals $\mathscr{I}_{X,\xi}$ and $\mathscr{I}_{Y,\xi}$ pull back to invertible ideals supported on the pullbacks of those $E_{i}$ meeting $\mu^{-1}(\xi)$.
@@
For (2), fix $\xi\in D_{X}\cap D_{Y}$ and set $Z_{\xi}:=Z\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$. Since $D_{X}$ and $D_{Y}$ are Cartier divisors on $\overline{\Gamma}_{F}$, their ideal sheaves $\mathscr{I}_{X}$ and $\mathscr{I}_{Y}$ are invertible. Hence, their pullbacks to $Z_{\xi}$ remain invertible ideal sheaves and define Cartier divisors supported on the pullback of $(D_{X})_{\mathrm{red}}+(D_{Y})_{\mathrm{red}}$.
By (1), the divisor $(D_{X})_{\mathrm{red}}+(D_{Y})_{\mathrm{red}}$ is simple normal crossings on $Z$, with irreducible components $E_{1},\dots,E_{N}$. Thus, locally on $Z$, it is defined by an equation $z_{1}\cdots z_{r}=0$, where $z_{j}=0$ cuts out one of the components $E_{i}$. After base change to $\operatorname{Spec}(R_{\xi})$, the pullback divisor is still locally defined by $z_{1}\cdots z_{r}=0$, so its irreducible components are exactly the pullbacks $E_{i}\times_{\overline{\Gamma}_{F}}\operatorname{Spec}(R_{\xi})$ for those $i$ with $E_{i}\cap\mu^{-1}(\xi)\neq\varnothing$. Therefore the pullbacks of $\mathscr{I}_{X}$ and $\mathscr{I}_{Y}$ to $Z_{\xi}$ are invertible ideal sheaves whose supports are contained in those components. ∎
@@ 12.2. Uniform verification on a fixed toric model
$\mathfrak{a}_{t}\ =\ \pi_{*}\mathcal{O}_{Y}\Big(-\sum_{i=1}^{N}d_{i}(t)E_{i}\Big),$
where $d_{i}(t)=\sum_{j=1}^{r}a_{ij}\lceil\lambda_{j}t\rceil$. In particular, for any other family $\mathfrak{b}_{\bullet}$} of the same form (possibly with different coefficients), the containment relations $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ are determined by the finitely many divisorial valuations $\{\operatorname{\mathrm{ord}}_{E_{i}}\}_{i=1}^{N}$, i.e., the hypotheses of the finite-testing criterion in Theorem 5.8 are satisfied with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$.
@@
$D_{\mathfrak{a}_{\bullet}}(t)\ :=\ \sum_{j=1}^{r}\lceil\lambda_{j}t\rceil\,F_{j}\ =\ \sum_{i=1}^{N}d_{i}(t)E_{i},\text{ with }d_{i}(t):=\sum_{j=1}^{r}a_{ij}\lceil\lambda_{j}t\rceil\in\mathbb{N}.$
@@
$\Big(\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}\Big)\mathcal{O}_{Y}\ =\ \mathcal{O}_{Y}\big(-D_{\mathfrak{a}_{\bullet}}(t)\big).$
@@
$\mathfrak{a}_{t}\ =\ \overline{\prod_{j=1}^{r}(I^{(j)})^{\lceil\lambda_{j}t\rceil}}\ =\ \pi_{*}\mathcal{O}_{Y}\big(-D_{\mathfrak{a}_{\bullet}}(t)\big).$
Step 2: finite testing. For any $p,q$, the condition $\mathfrak{b}_{q}\subseteq\mathfrak{a}_{p}$ is equivalent to the inequality of divisors on $Y$: $D_{\mathfrak{b}_{\bullet}}(q)\ \geq\ D_{\mathfrak{a}_{\bullet}}(p)$, i.e., coefficientwise inequalities along $E_{i}$. Equivalently, $\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{b}_{q})\geq\operatorname{\mathrm{ord}}_{E_{i}}(\mathfrak{a}_{p})$ for all $i$. This is precisely the finite-testing hypothesis in Theorem 5.8 with testing set $\{\operatorname{\mathrm{ord}}_{E_{i}}\}$. ∎